# Pangenome-guided sequence assembly via binary optimisation

### v1

@@ Abstract
De novo genome assembly is challenging in highly repetitive regions; however, reference-guided assemblers often suffer from bias. We propose a framework for pangenome-guided sequence assembly, which can resolve short-read data in complex regions without bias towards a single reference genome. Our method frames assembly as a graph traversal optimisation problem, which can be implemented on quantum computers. The pipeline first annotates pangenome graphs with estimated copy numbers for each node, then finds a path on the graph that best explains those copy numbers. On simulated data, our approach significantly reduces the number of contigs compared to de novo assemblers. While it introduces a small increase in inaccuracies, such as false joins, our optimisation-based methods are competitive with current exhaustive search techniques. They are also designed to scale more efficiently as the problem size grows and will run effectively on future quantum computers.
@@ 1 Introduction
A single linear reference genome cannot capture the genome variation that occurs within species, whether the variation is local, such as Single Nucleotide Variants (SNVs), or larger structural changes, for example Copy Number Variations (CNVs) [1]. Bioinformatic pipelines often include a step where a new sample is compared to a reference; for example, in variant detection. When a single reference is used, reference bias is inevitable: a tendency to report fragments of sequence that appear in the reference, and to miss novel or highly divergent regions [1, 2, 3]. The solution is to use references that capture genomic diversity; the study of such objects is pangenomics.
@@
Here we propose a methodology for solving pangenome-guided sequence assembly by directly mapping reads onto a pangenome, then recasting assembly as a graph traversal problem. In most of the genome, traversal is trivial. We focus on highly complex regions of the pangenome graph. A sketch of our pipeline is shown in fig. 1, and consists of 4 steps:
  1. 1.
    Short reads are aligned onto nodes of a pangenome.
  2. 2.
    Copy numbers for each node are estimated from the coverage data.
  3. 3.
    A path through the pangenome graph that best explains the copy numbers is found.
  4. 4.
    The corresponding sequence is extracted and optionally post-processed to improve assembly quality.
Figure 1: A sketch of the pangenome-guided sequence assembly procedure. (Upper-left) A pangenome graph and a set of short reads sequenced from a new individual. (Upper-right) Reads are aligned onto nodes of the graph. (Lower-left) Copy numbers are estimated from the coverage of each node. (Lower-right) A path that best explains the data is found, and the corresponding sequence is extracted.
@@
The key computational task in the pipeline is the path-finding in step 3. This problem is similar to those tackled by certain de novo genome assemblers like Flye [12] and Oatk [13], which aim to resolve repeat-induced non-linear structures in assembly graphs. Their approach typically involves estimating the copy numbers of repeat sequences from sequence coverage and then identifying a graph traversal that satisfies these copy number constraints. In the context of pangenome graphs, however, additional complexities arise. First, due to genomic divergence, a pangenome graph does not fully represent the target genome, necessitating imputation of missing or divergent regions. Second, read mapping to a pangenome graph is inherently noisier, largely due to its repetitive and multi-genome nature, resulting in greater uncertainty in sequence coverage and, consequently, in copy number estimates. To accurately resolve the graph paths through repetitive regions, pangenome-based methods must be more robust to such noise.
[…]
@@ 2 Results
Our main comparison point for existing state-of-the-art is pathfinder [13]. Pathfinder is a submodule from the oat-k package, a tool for resolving plant organelles using high-accuracy long reads. While this application is far from our current setting, the methods are somewhat related. In particular, pathfinder also assigns weights to a graph and then searches for a walk that explains the data. However, it proceeds via an exhaustive search rather than an optimisation. To the best of our knowledge, this is the software package that gives the fairest comparison with our results.
@@
A comparison of pangenome-guided assembly of synthetic haploid data via kmer mapping using pathfinder against de novo sequence assembly can be seen in table 1. The primary difference lies in the number of contigs returned and demonstrates the great benefit of using a pangenome-guided approach to short-read assembly. However, this does come with a small increase in false joins (seen as breaks in alignment against the truth set). Optimising the path sequence by realigning the short read data against it and producing a new consensus does improve accuracy across a range of measures. The percentage identity remains slightly lower than for the de novo assemblers, indicating a hybrid solution may be the optimal strategy.
@@
Annotator/Assembler %Covered %Used Contigs Breaks Indel No.Diff %Identity
Kmer2node 86.7 93.2 2.8 1.9 1.3 0.2 96.9
Minigraph 84.0 90.9 4.7 2.4 1.0 0.0 97.2
GraphAligner 92.4 91.8 1.7 2.0 1.6 0.2 96.9
Giraffe 91.8 91.5 1.6 2.3 1.5 0.1 96.9
Kmer2node-opt 86.8 94.1 2.6 1.5 0.6 0.3 98.7
Minigraph-opt 84.4 92.7 4.3 1.8 0.3 0.1 99.2
GraphAligner-opt 92.7 92.5 1.6 1.7 0.8 0.3 98.6
Giraffe-opt 92.1 92.0 1.5 1.9 0.8 0.3 98.6
Syncasm 91.6 86.2 31.3 0.0 0.1 0.0 100.0
Miniasm 81.5 91.3 14.8 0.1 0.1 0.0 99.6
@@
Table 1: Performance of pangenome guided mapping of short-read data using minigraph, GraphAligner, vg giraffe and kmer2node weight assignments with the pathfinder solver, compared against pure de novo sequence assembly from syncasm and miniasm. Averaged stats from 51 seeds with 5 alignments each. Pathfinder results are presented twice, evaluating the original concatenated path and an optimised path formed by realigning all the short-read data back to the original path to produce a refined consensus sequence.
We test the optimisation formulation across a range of solvers. These employ different amounts of “quantumness”, ranging from purely classical solvers, through quantum annealing (QA), which is tailored specifically for quantum optimisation, to the full power of quantum circuits. All quantum solvers are used as part of a hybrid classical-quantum algorithm, where a quantum processor performs the “hard core” of the algorithm and a classical computer handles simpler tasks such as pre-processing and data storage.
@@ 2.1 Classical optimisers
We test each combination on a set of 20 pangenomes with 5 sequences to be aligned against each pangenome. The pangenomes have an average of 77.8 nodes. Gurobi and MQLib are given time limits of 5 and 300 seconds to test whether higher quality solutions are obtained at long run times. These classical optimisers are allowed 3 runs per sequence and per time limit to account for variations in solution quality due to their heuristic nature. The averaged results are given in table 2. Radar charts showing the average performance across the metrics discussed in section 4.4 are given in fig. 2. Figure 2(a) shows that, for any choice of annotation strategy, each of the solvers is broadly competitive with the others. When either kmer2node or minigraph is used, the average case results across each of the solvers are very similar. The starkest difference comes when GraphAligner is used. In that case, pathfinder reports more large differences than either MQLib or Gurobi, but fewer breaks and slightly fewer large indels on average. The large number of breaks in the QUBO solver results may be due to not taking into account edge weights, which can help to resolve inversions: see section 3 for a brief discussion.
Annotator Solver %Cov. %Used Contigs Breaks Indels Diffs %Identity
GraphAligner Gurobi 85.69 83.85 1.92 5.31 1.09 0.18 98.45
MQLib 86.90 77.75 1.73 6.76 1.08 0.19 98.42
Pathfinder 86.74 91.42 1.73 2.09 0.97 0.33 98.21
Kmer2node Gurobi 82.50 92.07 3.73 2.17 0.86 0.17 98.82
MQLib 82.36 92.88 3.68 1.95 0.90 0.15 98.75
Pathfinder 81.55 92.01 3.85 1.76 0.88 0.19 98.70
Minigraph Gurobi 80.09 89.92 6.74 1.70 0.63 0.12 99.09
MQLib 79.40 90.31 6.47 1.52 0.56 0.09 99.14
Pathfinder 78.51 90.06 6.74 1.52 0.52 0.11 99.17
@@
Table 2: Performance of pangenome guided mapping of short-read data using minigraph, GraphAligner and kmer2node weight assignments with the MQLib, Gurobi and pathfinder solvers. Averaged stats from 20 seeds with 5 alignments each. MQLib and Gurobi were given 300 seconds per run, and 3 runs per alignment.
@@
(a) Radar charts comparing the performance of each of the annotation strategies.
(b) Radar charts comparing the performance of each of the classical solvers.
(c) A violin plot comparing the performance of each combination of annotation strategy and classical solver.
@@
Figure 2: Radar charts comparing the performance of each combination of annotation strategy: GraphAligner, kmer2node and minigraph; and classical solver: Gurobi, MQLib and pathfinder. The seven axes plotted are the evaluation criteria discussed in section $4.$ For each, the number in brackets corresponds to the outermost point on that axis. For the lower half of each plot, further out along the axis corresponds to fewer contigs, breaks and so on. (c) Violin plots showing the per-instance performance of Gurobi, MQLib and pathfinder.
@@
As well as average-case performance, it is illuminating to see how the combinations perform on specific instances, particularly hard ones. Violin plots showing the average of the “Covered” and “Used” statistics across all instances are given in fig. 2(c). When GraphAligner is used to annotate graphs, it is clear that pathfinder outputs high-quality solutions on a majority of instances, although a very low worst-case outcome drags down its average performance. For the other annotators, the QUBO solvers are much more competitive, with similarly shaped plots across both the minigraph and kmer2node settings.
@@ 2.2 Quantum annealing
To test the potential of QA, we perform a range of experiments using the Leap hybrid solver provided by the quantum hardware company D-Wave [21]. This solver is provided on the cloud, and typically one or more classical algorithms run on the problem while hard parts are outsourced to the QPU. A configurable time limit is provided, and the best solution found during that time is returned, along with some metadata. The main benefit of QA is the size of the hardware available. The Advantage2 system has 4,400 qubits arranged in a lattice structure, around 30 times larger than superconducting offerings from IBM. We use the QA systems to wayfind the potential for quantum advantage on more standard quantum hardware in the future.
@@
Annotator Solver %Cov. %Used Contigs Breaks Indels Diffs %Identity
GraphAligner D-Wave 80.64 86.31 1.10 5.33 0.63 0.10 99.03
pathfinder 93.24 92.29 1.10 1.77 0.83 0.10 99.22
kmer2node D-Wave 85.02 92.45 1.97 2.87 0.57 0.20 98.87
pathfinder 85.69 94.44 1.90 1.27 0.77 0.13 99.08
minigraph D-Wave 86.29 94.29 3.50 1.63 0.23 0.03 99.59
pathfinder 86.03 94.43 3.53 1.27 0.33 0.03 99.58
Table 3: Performance of pangenome guided mapping of short-read data using minigraph, GraphAligner and kmer2node weight assignments with the D-Wave and pathfinder solvers. Averaged stats from 10 seeds with 3 alignments each. D-Wave was given 60 seconds per run, and 2 runs per alignment.
(a) Radar charts comparing the performance of each of the annotation strategies.
(b) Radar charts comparing the performance of D-Wave and pathfinder.
(c) A violin plot comparing the performance of D-Wave and pathfinder for each annotation strategy.
@@
2.3 Classical simulations of quantum circuits
We also test hybrid classical-quantum optimisation algorithms, with the quantum part being simulated classically. While the classical simulation limits the size of the problem instance we can explore, our results demonstrate the feasibility of our method for future runs on real hardware. We use various techniques to improve the convergence of the algorithm – see section 4.8 for details. Figure 4 shows the results of simulating the QAOA algorithm with $p=2$ solving a small Oriented Tangle Resolution problem. The circuit has 28 qubits and 1908 2-qubit gates after being compiled to the IBM-native gate set, with a 2-qubit depth of 146. 1000 samples were simulated at each iteration, and a maximum of 100 iterations were allowed. We see that the energy has effectively converged after around 25 iterations, and the classical optimiser terminates after 78 iterations. The total simulation took under 4 minutes to run on a single GPU. Figure 4(b) demonstrates the successful convergence of the QAOA procedure, with the vast majority of QAOA samples lying at a lower objective value than the best random samples.
(a) The objective value at each iteration of a QAOA simulation.
(b) Histogram of the samples obtained at the final iteration of the QAOA algorithm (blue), compared with random samples (orange). The optimal solution is shown in black dashes.
Figure 4: Plots showing the performance of a 28-qubit QAOA simulation for Oriented Tangle Resolution.

### v2

@@ Abstract
De novo genome assembly is challenging in highly repetitive regions; however, reference-guided assemblers often suffer from bias. We propose a framework for pangenome-guided sequence assembly, which can resolve short-read data in complex regions without bias towards a single reference genome. Our primary contribution is to frame the assembly as a graph traversal optimisation problem, which can be implemented classically or on a quantum computer. The workflow involves first annotating pangenome graphs with estimated copy numbers for each node, then finding a path on the graph that best explains those copy numbers. On simulated data, our approach significantly reduces the number of contigs compared to de novo assemblers. While they introduce a small increase in inaccuracies, such as false joins, our optimisation-based methods are competitive with current exhaustive search techniques. They are also designed to scale more efficiently as the problem size grows and will run effectively on future quantum computers; a small experiment on a real quantum device showcases this behaviour. Moreover, they are more resilient to noise in copy number estimation inherent in short-read-based assembly. We also develop novel tools for creating realistic synthetic pangenomes, aligning reads to pangenomes and for evaluating assembly quality.
Keywords: pangenomics, assembly, optimisation, quantum algorithms
Authors: all authors are part of a research collaboration participating in the Wellcome Leap Q4Bio programme, of which Sergii Strelchuk is the main PI and Richard Durbin a co-PI.
@@ 1 Introduction
A single linear reference genome alone cannot capture the genome variation that occurs within species, whether the variation is local, such as Single Nucleotide Variants (SNVs), or larger structural changes, like Copy Number Variations (CNVs) [3]. Bioinformatic pipelines often include a step where a new sample is compared to this fixed reference genome; for example, in variant detection. When a single reference is used, reference bias is inevitable: a tendency to report fragments of sequence that appear in the reference, and to miss novel or highly divergent regions [3, 17, 12]. The solution is to use references that capture genomic diversity; one such object is the pangenome.
@@
Here, we propose a methodology for pangenome-guided sequence assembly that maps reads onto a pangenome and reformulates the assembly as a graph-traversal optimisation. Most genomic regions are easily resolved; we focus on the highly complex, repetitive regions of the pangenome graph. The pipeline, summarised in fig. 1, consists of five stages:
1. Problem creation:
Generate a synthetic pangenome and a new individual genome. Simulate short-read sequencing (typically 30× coverage).
2. Read mapping:
Align short reads to the pangenome using GraphAligner, kmer2node or minigraph, and annotate nodes with kmer counts.
3. Copy number estimation:
Estimate node copy numbers from the annotated graph.
4. Path finding:
Identify a path through the pangenome graph that best fits the observed copy numbers using either the pathfinder tool or a novel binary optimisation approach.
5. Solution processing:
Extract and optionally re-align the resulting sequence to improve quality before evaluating against the ground truth.
A flow chart showing the steps involved in pangenome-guided sequence assembly.
Figure 1: A sketch of the pangenome-guided sequence assembly procedure. (1) Problem creation consists of two steps. A pangenome and a new individual genome are synthesised, and the genome is shredded to simulate shotgun sequencing. (2) Read mapping involves aligning the short reads to the pangenome using 1 of 3 software tools and tagging the nodes with the observed kmer counts. (3) Copy number estimation is performed using pathfinder. (4) Optimal path finding depends on the choice of solver. If using pathfinder, input the annotated graph directly. Otherwise, construct the QUBO matrix from the graph, and input that into the chosen QUBO solver. (5) Solution processing starts with an optional re-alignment step to improve solution quality, before evaluating solution quality with a variety of metrics.
@@ 5. Solution processing:
The key computational task in the pipeline is the path finding in step 4. This problem is similar to those tackled by certain de novo genome assemblers like Flye [19] and Oatk [33], which aim to resolve repeat-induced non-linear structures in assembly graphs. Their approach typically involves estimating the copy numbers of repeat sequences from sequence coverage and then identifying a graph traversal that satisfies these copy number constraints. They are designed for datasets derived from a single individual, in which the resulting graphs are typically complete and relatively noise-free, allowing the estimation of sequence copy numbers reliably.
In the context of pangenome graphs, however, additional complexities arise. First, due to genomic divergence, a pangenome graph does not fully represent the target genome, necessitating imputation of missing or divergent regions. Second, read mapping to a pangenome graph is inherently noisier due to its repetitive and multi-genome nature, resulting in greater uncertainty in sequence coverage and, consequently, in copy number estimates. To accurately resolve the graph paths through repetitive regions, pangenome-based methods must be more robust to such noise.
To that end, we formalise the path finding problem as an optimisation problem. The optimisation may be viewed as a generalisation of the Hamiltonian path problem and is $\mathsf{NP}$}-hard in general. We cast it as a Quadratic Unconstrained Binary Optimisation (QUBO), a framework commonly used in the optimisation literature due to its ability to strike a balance between expressivity and tractability. Moreover, QUBO is in some sense “native” for quantum optimisation, particularly the Quantum Adiabatic Optimisation Algorithm (QAOA). Quantum computers are currently limited in size and error-prone, characteristic of the so-called Noisy Intermediate-Scale Quantum (NISQ) era. In the NISQ era, applications of quantum computers must be carefully chosen to mitigate these limitations; combinatorial optimisation is a promising direction that has already seen some success [18, 24, 26, 28]. A primer on quantum computing and quantum optimisation is given in section S1. Below, we compare both classical and quantum approaches to the optimisation and discuss the potential utility of quantum approaches.
@@ 2 Results
Our main comparison point for existing state-of-the-art is pathfinder [33]. Pathfinder is a submodule from the Oatk package, a tool for resolving plant organelles using high-accuracy long reads. While this application is far from our current setting, the methods are somewhat related. In particular, pathfinder also assigns weights to a graph and then searches for a walk that explains the data. However, it proceeds via exhaustive search rather than optimisation; moreover, pathfinder was designed to resolve assembly graphs whose annotated coverage values are essentially noiseless and have no unused nodes. Conversely, in our setting, the reads originate from an individual not present in the pangenome, and we therefore expect imperfect coverage values and many nodes to have zero coverage. We use a cost function to be robust to this noise.
As an example, consider a simple line graph with nodes $A\rightarrow B\rightarrow C\rightarrow D\rightarrow E$ and estimated copy numbers (1, 1, 0, 1, 1)(1,\,1,\,0,\,1,\,1). The prescriptive nature of pathfinder yields two contigs: $A\rightarrow B$ and $C\rightarrow D$. Conversely, the cost function used in our QUBO formulation allows a single unified path of $A\rightarrow\ldots\rightarrow E$ to be found. Nevertheless, to the best of our knowledge, pathfinder is the software package that gives the fairest comparison with our results.
@@
A comparison of pangenome-guided assembly of synthetic haploid data via kmer mapping using pathfinder against de novo sequence assembly can be seen in table 1. The primary difference lies in the number of contigs returned and demonstrates the great benefit of using a pangenome-guided approach to short-read assembly. However, this comes with a small increase in false joins (i.e., breaks in alignment with the truth set). Optimising the path sequence by realigning the short-read data against it and producing a new consensus improves accuracy across a range of metrics. The percentage identity remains slightly lower than that of the de novo assemblers, indicating that a hybrid solution may be the optimal strategy.
@@
Annotator/Assembler %Covered %Used Contigs Breaks Indel No.Diff N50 %Identity
Kmer2node 88.3 93.2 2.8 1.9 1.2 0.2 6697 96.9
Minigraph 85.1 91.2 4.7 2.3 1.0 0.0 5044 97.2
GraphAligner 92.6 92.0 1.7 2.0 1.7 0.2 9484 96.9
Giraffe 91.7 91.7 1.9 2.0 1.4 0.1 8748 97.0
Kmer2node-opt 88.4 94.0 2.6 1.5 0.6 0.3 6679 98.7
Minigraph-opt 85.4 92.9 4.3 1.8 0.3 0.1 5002 99.2
GraphAligner-opt 93.0 92.6 1.6 1.6 0.8 0.3 9436 98.6
Giraffe-opt 91.9 92.6 1.8 1.7 0.7 0.3 8701 98.7
Syncasm 91.6 86.2 31.3 0.0 0.1 0.0 878 100.0
Miniasm 81.5 91.3 14.8 0.1 0.1 0.0 1022 99.6
@@
Table 1: Performance of pangenome guided mapping of short-read data using minigraph, GraphAligner, vg giraffe and kmer2node weight assignments with the pathfinder solver, compared against pure de novo sequence assembly from syncasm and miniasm. Weight assignment methods are discussed in detail in section S3. Averaged stats from 51 seeds with 5 alignments each. Pathfinder results are presented twice, evaluating the original concatenated path and an optimised path formed by realigning all the short-read data back to the original path to produce a refined consensus sequence. Columns represent the proportion of the true genome covered by the assembly result (%Covered), and the percentage of this result covered by the true genome (%Used), the number of contigs, the number of breaks in alignment of assembly to true sequence (breaks, i.e. false joins), the number of large indels and regions of significant runs of differences, the N50 contig size, and total percent identity between assembly and true genome. See section 4.4 for further details.
To evaluate the optimisation formulation, we tested a range of solvers with varying levels of quantum integration. These methods span a spectrum from traditional classical solvers, to hybrid approaches like quantum annealing QA, which is specialised for optimisation problems, to gate-based quantum computers.
@@ 2.1 Classical optimisers
We test each combination on a set of 20 pangenomes with 5 sequences to be aligned against each pangenome. The pangenomes have an average of 77.8 nodes. Gurobi and MQLib are given time limits of 5 and 300 seconds to test whether higher quality solutions are obtained at long run times. These classical optimisers are allowed 3 runs per sequence and per time limit to account for variations in solution quality due to their heuristic nature. The averaged results are given in table 2. Radar charts showing the average performance across the metrics discussed in section 4.4 are given in fig. 2. Figure 2a shows that, for any choice of annotation strategy, each of the solvers is broadly competitive with the others. When either kmer2node or minigraph is used, the average case results across each of the solvers are very similar. The starkest difference comes when GraphAligner is used. In that case, pathfinder reports more large differences than either MQLib or Gurobi, but fewer breaks and slightly fewer large indels on average. The large number of breaks in the QUBO solver results may be due to not taking into account edge weights, which can help to resolve inversions: see section 3 for a brief discussion.
Annotator Solver %Cov. %Used Contigs N50
GraphAligner Gurobi 85.69 (12.46) 83.85 (11.75) 1.92 (1.52) 9755 (2699)
MQLib 86.90 (12.97) 77.75 (13.60) 1.73 (1.13) 11337 (3846)
Pathfinder 86.74 (15.82) 91.42 (8.56) 1.73 (1.13) 8725 (2194)
Kmer2node Gurobi 82.50 (12.52) 92.07 (7.68) 3.73 (2.50) 5370 (2529)
MQLib 82.36 (12.19) 92.88 (6.44) 3.68 (2.27) 5201 (2523)
Pathfinder 81.55 (13.42) 92.01 (7.76) 3.85 (2.71) 5208 (2586)
Minigraph Gurobi 80.09 (11.74) 89.92 (7.37) 6.74 (3.69) 3512 (2708)
MQLib 79.40 (13.45) 90.31 (7.59) 6.47 (3.58) 3468 (2696)
Pathfinder 78.51 (13.84) 90.06 (7.48) 6.74 (3.84) 3409 (2684)
Annotator Solver Breaks Indels Diffs %Identity
GraphAligner Gurobi 5.31 (3.76) 1.09 (1.09) 0.18 (0.48) 98.45 (1.38)
MQLib 6.76 (4.90) 1.08 (1.14) 0.19 (0.44) 98.42 (1.35)
Pathfinder 2.09 (2.08) 0.97 (1.18) 0.33 (0.62) 98.21 (1.90)
Kmer2node Gurobi 2.17 (2.09) 0.86 (1.04) 0.17 (0.40) 98.82 (1.14)
MQLib 1.95 (1.72) 0.90 (1.14) 0.15 (0.36) 98.75 (1.32)
Pathfinder 1.76 (1.53) 0.88 (1.05) 0.19 (0.42) 98.70 (1.34)
Minigraph Gurobi 1.70 (1.58) 0.63 (0.95) 0.12 (0.35) 99.09 (0.93)
MQLib 1.52 (1.40) 0.56 (0.82) 0.09 (0.29) 99.14 (0.75)
Pathfinder 1.52 (1.34) 0.52 (0.79) 0.11 (0.34) 99.17 (0.91)
@@
Table 2: Performance of pangenome guided mapping of short-read data using minigraph, GraphAligner and kmer2node weight assignments with the MQLib, Gurobi and pathfinder solvers. Averaged stats from 20 seeds with 5 alignments each, with standard deviations given in parentheses. MQLib and Gurobi were given 300 seconds per run, and 3 runs per alignment.
@@
A set of radar charts and violin plots comparing the relative performance of combinations of annotation strategies and classical solvers.
@@
Figure 2: (a), (b) Radar charts comparing the performance of each combination of annotation strategy: GraphAligner, kmer2node and minigraph; and classical solver: Gurobi, MQLib and pathfinder. The seven axes plotted are the evaluation criteria discussed in section $4.$ For each, the number in brackets corresponds to the outermost point on that axis. For the lower half of each plot, further out along the axis corresponds to fewer contigs, breaks and so on. (c) Violin plots showing the per-instance performance of Gurobi, MQLib and pathfinder.
@@
As well as average-case performance, it is illuminating to see how the combinations perform on specific instances, particularly hard ones. Violin plots showing the average of the “Covered” and “Used” statistics across all instances are given in Figure 2c. When GraphAligner is used to annotate graphs, it is clear that pathfinder outputs high-quality solutions on a majority of instances, although a very low worst-case outcome drags down its average performance. For the other annotators, the QUBO solvers are much more competitive, with similarly shaped plots across both the minigraph and kmer2node settings.
@@ 2.2 Quantum annealing
To test the potential of QA, we perform a range of experiments using the Leap hybrid solver provided by the quantum hardware company D-Wave [8]. This solver is provided on the cloud, and typically one or more classical algorithms run on the problem while hard parts are outsourced to the QPU. A configurable time limit is provided, and the best solution found during that time is returned, along with some metadata. The main benefit of QA is the size of the hardware available. The Advantage2 system has 4,400 qubits arranged in a lattice structure, around 30 times larger than superconducting offerings from IBM and around 90 times larger than Quantinuum’s trapped ion computers. We use the QA systems to wayfind the potential for quantum advantage on more standard quantum hardware in the future.
@@
Annotator Solver %Cov. %Used Contigs N50
GraphAligner D-Wave 80.18 (13.25) 85.07 (13.49) 1.10 (0.54) 9062 (1727)
pathfinder 93.24 (11.76) 92.29 (11.00) 1.10 (0.54) 10060 (385)
kmer2node D-Wave 85.45 (11.31) 91.52 1(8.55) 1.97 (1.60) 7203 (2315)
pathfinder 85.99 (16.44) 93.99 1(8.65) 2.00 (1.59) 7330 (2586)
minigraph D-Wave 85.99 (10.85) 94.75 1(6.37) 3.53 (1.71) 4361 (1879)
pathfinder 86.03 (11.15) 94.43 1(6.93) 3.53 (1.71) 4492 (1913)
Annotator Solver Breaks Indels Diffs %Identity
GraphAligner D-Wave 5.77 (4.26) 0.53 (0.96) 0.07 (0.25) 99.18 (0.98)
pathfinder 1.77 (2.73) 0.83 (1.00) 0.10 (0.30) 99.22 (1.15)
kmer2node D-Wave 2.97 (2.34) 0.50 (0.76) 0.13 (0.34) 99.08 (1.13)
pathfinder 1.23 (1.52) 0.70 (0.90) 0.13 (0.34) 99.14 (1.03)
minigraph D-Wave 1.30 (1.07) 0.23 (0.50) 0.03 (0.18) 99.55 (0.53)
pathfinder 1.27 (1.26) 0.33 (0.70) 0.03 (0.18) 99.58 (0.53)
Table 3: Performance of pangenome guided mapping of short-read data using minigraph, GraphAligner and kmer2node weight assignments with the D-Wave and pathfinder solvers. Averaged stats from 10 seeds with 3 alignments each, with standard deviations given in parentheses. D-Wave was given 60 seconds per run, and 2 runs per alignment.
A set of radar charts and violin plots comparing the relative performance of combinations of annotation strategies and either the D-Wave or pathfinder solvers.
@@
2.3 Quantum optimisation
We also test hybrid classical-quantum optimisation algorithms as a proof-of-concept. We use various techniques to improve the convergence of the algorithm – see section 4.8 for details. For the purposes of testing our algorithms, the majority of runs are performed with the quantum computer being simulated classically. We can simulate circuits up to 35 qubits using heterogeneous CPU + GPU simulations. However, we also perform small experiments on real quantum devices; a comparison of simulated and actual performance for one such experiment is provided below.
Figure 4 shows the results of simulating the QAOA algorithm with $p=4$ on a small Oriented Tangle Resolution problem. The circuit has 10 qubits, 496 2-qubit gates and a 2-qubit depth of 112 after being compiled to the IBM-native gate set. The chosen emulator implements noiseless gates, but we include sampling noise: 256 samples are taken at each iteration. A maximum of 100 iterations were allowed. The simulation took under 20 seconds to run on a single CPU and GPU, including circuit compilation time. Figure 4 demonstrates the successful convergence of the QAOA procedure. The simulated quantum computer sampled the optimum repeatedly and had several other samples at the same energy as the best random samples.
Figure 5 shows the results of running the same QAOA algorithm, with the same initial conditions, on an actual quantum device. The device used is the IBM Strasbourg computer, an Eagle R3 device with 127 qubits in a “heavy-hex“ layout and an average 2-qubit error rate of around $3\times 10^{-}2$. The full computation took 213 seconds. Each of the 55 iterations of the algorithm used less than 1s of QPU time; the remaining time is comprised of classical computation, communication with the device and preparing the device for the next iteration. Clearly, noise in the device has degraded the convergence of the classical optimiser. Nonetheless, Figure 5b still shows a non-trivial shift of probability mass towards lower-energy solutions. Moreover, the optimal solution was again sampled, and several other low-energy solutions were found.
Going forward, it is clear that error mitigation techniques are critical for the QAOA algorithm to succeed. Since we are essentially considering a sampling task, measurement error mitigation will likely be a good choice. The implementation of these techniques is left for future work.
A pair of plots showing the convergence of the energy during a simulated QAOA run, and the samples drawn at the final iteration of the run.
Figure 4: Plots showing the performance of a 10-qubit QAOA simulation for Oriented Tangle Resolution.
A pair of plots showing the convergence of the energy during a hardware QAOA run, and the samples drawn at the final iteration of the run.
Figure 5: Plots showing the performance of a 10-qubit QAOA simulation for Oriented Tangle Resolution.